utility function is u(x,y) = 6x^.5+y , and his budget constraint is px*x +y = m,

m=30

px=2

py = 1.

a.What is his original highest utility

level?

Now px has decreased to 1, m and py do not change.

b.What is his new maximum utility level?


1
Expert's answer
2021-10-28T08:49:29-0400

Solution:

a.). Utility function (UX,Y) = 6x0.5 + y


To maximize utility: "\frac{MU_{X} }{P_{X}} = \frac{MU_{Y} }{P_{Y}}"


MUX = "\frac{\partial U} {\partial X}" = 3x-0.5


MUY = "\frac{\partial U} {\partial Y} = 1"


Price of x = 2

Price of y = 1


"\frac{3X^{-0.5} }{2} = \frac{1}{1}"


X = 2.25

Budget constraint: PxX + Y = M

2X + Y = 30

X = 2.25

2(2.25) + Y = 30

4.5 + Y = 30

Y = 25.5

Highest utility: U(x,y) = (2.5, 25.5)

 

b.). New budget constraint: PxX + Y = M

X + Y = 30

X = 2.25

2.25 + Y = 30

Y = 27.75

New maximum utility: U(x,y) = (2.5, 27.75)